What is the value of the angle, if 100 degrees less than two times the supplement of the angle is thrice of its complement?
Answer by ewatrrr(24785)

 
Hi
Note: Sum of an angle and it's supplement = 180°
Let x and (180-x)represent the measure of the complement and it's supplement
Question states***
100 degrees less than two times the supplement of the angle is thrice of its complement
Write as we Read:
2(180-x)-100 = 3x
360 - 2x - 100=3x
260= 5x
52° = x is the complement, it's supplement is 128°
CHECKING our Answer***
256-100=156 = 3· 52